OpenStudy (anonymous):

A car dealership has 24 sedans and 18 sports coupes on its lot. What is the ratio of sedans to sports coupes on the lot? A. 3 : 4 B. 3 : 2 C. 2 : 3 D. 4 : 3

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

24 sedans to 18 sports coupes to indicates : so 24:18

OpenStudy (hockeychick23):

4:3

OpenStudy (anonymous):

then simplify

Mehek (mehek14):

simplify that and you get 4:3

OpenStudy (anonymous):

divide them both by 3, 24 / 3 = 8 18 / 3 = 6 8 : 6 is not an answer, so divide by 2, 8 /2 = 4 6 / 2 = 3 4 : 3

OpenStudy (anonymous):

Stephanie made 36 muffins. She made 24 apple muffins and the rest were banana. What is the ratio of apple muffins to banana muffins? A. 2 to 3 B. 12 to 36 C. 2 to 1 D. 24 to 36

OpenStudy (anonymous):

Which ratio correctly compares 75 cm to 60 mm? A. 5 : 4 B. 3 : 2 C. 75 : 60 D. 25 : 2

OpenStudy (anonymous):

divide both by 6 36 / 6 = 6 24 / 6 = 4 divide them both by 2 6 / 2 = 3 4 / 2 = 2 3 : 2

OpenStudy (anonymous):

one second, for the second question

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

divide both by 5 75 / 5 = 15 60 / 5 = 12 divide them both by 3 15 / 3 = 5 12 / 3 = 4 5 : 4

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

medal please ?